### 13 Feb 2024 v2

# Discriminative Adversarial Unlearning

Abstract
We introduce a novel machine unlearning framework founded upon the established principles of the min-max optimization paradigm. We capitalize on the capabilities of strong Membership Inference Attacks (MIA) to facilitate the unlearning of specific samples from a trained model. We consider the scenario of two networks, the attacker $\mathbf{A}$ and the trained defender $\mathbf{D}$ pitted against each other in an adversarial objective, wherein the attacker aims at teasing out the information of the data to be unlearned in order to infer membership, and the defender unlearns to defend the network against the attack, whilst preserving its general performance. The algorithm can be trained end-to-end using backpropagation, following the well known iterative min-max approach in updating the attacker and the defender. We additionally incorporate a self-supervised objective effectively addressing the feature space discrepancies between the forget set and the validation set, enhancing unlearning performance. Our proposed algorithm closely approximates the ideal benchmark of retraining from scratch for both random sample forgetting and class-wise forgetting schemes on standard machine-unlearning datasets. Specifically, on the class unlearning scheme, the method demonstrates near-optimal performance and comprehensively overcomes known methods over the random sample forgetting scheme across all metrics and multiple network pruning strategies. Code is available at: Link.
Machine Learning, ICML
1 Introduction
With over \$4 billion paid in settlements over privacy concerns by Big Tech firms since the enforcement of the General Data Protection Regulation (European Parliament and Council of the European Union, 2016) in 2018  [^1] Source: Data Privacy Manager, and more than 70% of citizens from nine major countries expressing support for increased governmental intervention in Big Tech [^2] Source: Amnesty International, significant concerns arise regarding individual privacy and the potential erosion of trust between human users and AI systems. Owing to their widespread incorporation, digital systems now function as vast repositories of personal data, creating a comprehensive digital footprint that reflects a broad spectrum of online behaviors, interactions, and communication patterns (Nguyên, 2019). This footprint is evident across various forms of user-generated content, including product reviews, blog posts, social media activities, and contributions to collaborative platforms such as Wikipedia (Nguyen et al., 2021; Carpentier et al., 2021). The application of artificial intelligence (AI) in areas such as targeted advertising and personalized product placement, along with the emergence of potentially harmful technologies like deepfakes, further underscores a critical concern regarding the lack of transparency and control over personal data, highlighting the need for protective measures in the management and utilization of personal information (Yevseiev et al., 2021; Al-Harrasi et al., 2023). The recent legal developments echo these concerns with a notable development being, ”the right to be forgotten,” (Dang, 2021; Tankard, 2016) which requires the deletion of personal data from both digital databases and machine learning models at an individual’s request. However, retraining AI models to accommodate each deletion request is impractical, given the potential cost of millions of dollars, coupled with substantial environmental impact.
In response to these challenges, there has been a notable surge in the development of algorithmic unlearning strategies (Wang et al., 2023; Zhang et al., 2023; Hoang et al., 2024; Chen et al., 2023; Lin et al., 2023; Liu et al., 2023a; Che et al., 2023). These strategies aim to modify existing models, allowing them to selectively ’unlearn’ specific samples from their training data. The approaches offer viable alternatives to complete retraining from scratch, providing cost-effective solutions that enhance data security and privacy while preserving the efficacy of the models. Despite their significant effectiveness, a majority of these measures (Liu et al., 2023a; Mehta et al., 2022; Gandikota et al., 2023; Heng and Soh, 2023a, b) primarily address the problem in the parametric space. They introduce modifications within the pre-training objectives (Lin et al., 2023) or the pre-processing phase (Liu et al., 2017; García-Pablos et al., 2020) of trained models, often entailing performance-cost trade-offs due to the evaluation of hessian approximations or proposals for complete overhauls of training procedures.
Membership Inference Attacks (MIA), on the other hand as a field of study, has concurrently bloomed (Shokri et al., 2017; Salem et al., 2018; Nasr et al., 2018; Leino and Fredrikson, 2020; Song and Mittal, 2021; Yuan and Zhang, 2022) exposing significant privacy related threats to trained models. For a trained model, MIAs are tasked with identification of inference of patterns in the model’s behaviour due to samples arising from its training set and other sets. Using these patterns, a successful attacker may avail themselves with personally identifying information (Veale et al., 2018; Fredrikson et al., 2015; Tramèr et al., 2016; Ganju et al., 2018) posing significant threats to the use of AI towards general welfare.
We posit that the objectives of the MIA and machine unlearning may be formulated as adversarially inclined and one may harness the recent developments in the fields in order to orchestrate a min-max framework and successfully conduct machine unlearning in pretrained models, preserving our trust in machine learning as a service (MLaaS). To our knowledge, our contribution is the first to formulate the problem of unlearning as an adversarial task wherein two networks, the attacker and the defender engage in a game against each other, which consequently leads towards efficient and effective unlearning measures learned in the defender, whilst notable retention of it’s original capacity for classification. We summarize our contribution in this work as follows.
  • We introduce a pioneering framework for machine unlearning, employing a min-max optimization procedure that engages neural networks in an end-to-end trainable game through backpropagation. Importantly, this framework eliminates the necessity for unrolling or approximation in the optimization procedures of the networks, and avoids the need for explicit gradient manipulation.
  • We innovate by repurposing the Barlow Twins objective (Zbontar et al., 2021) for unlearning, yielding a substantial performance boost. This repurposing opens avenues for applying self-supervised learning research to provide practical solutions in machine unlearning.
  • Through extensive experimentation, we showcase the effectiveness of our framework, achieving near-optimal performance and setting new benchmarks for machine unlearning on the standard machine-unlearning Cifar-10 and Cifar-100 datasets. Our framework excels under both random and class-wise forgetting scenarios.
2 Related Work
2.1 Machine Unlearning
Machine unlearning endeavors to mitigate the influence of specific data points or classes from a trained ML model, primarily addressing privacy concerns in accordance with regulations (Rosen, 2011; Hoofnagle et al., 2019). Recent developments in the field has led to a taxonomy for methods broadly classifying the objectives as follows.
Exact unlearning
Retraining the model from the scratch on the data filtering out forgetting dataset is the most straightforward and effective manner to update ML models for unlearning, ensuring complete deletion of undesirable samples or aspects from a pre-trained model. Cao and Yang (2015) explore this mode of unlearning over Naive Bayes classifiers whereas Ginart et al. (2019) implement deletion algorithms for $k\text{-}means$ clustering, rendering them unscalable to deep neural networks. Additionally, recent efforts (Bourtoule et al., 2021) also encourage training on shards of data, subsequently incorporating aggregation, effectively alleviating the expensive nature of training a large model from scratch. However, the measures signify impracticality in the face of large queries for deletion, requiring inordinate amounts of compute especially for models wherein the training cost estimates exceed millions of dollars.
Figure 1: Overview of the proposed framework. The figure depicts the interplay between the attacker and defender networks. The defender provides output and sensitivity information to the attacker which in turn provides feedback to the defender for unlearning. The objective is supplemented with a feature space self supervised regularization between the forget and validation sets.
Approximate unlearning
Acknowledgement of the said infeasibility of training large models from scratch has therefore led to an upsurge in palliative measures to selectively eradicate information from trained models. Model finetuning as an approach (Warnecke et al., 2021; Golatkar et al., 2020) aims to finetune the model on the retain set of data to induce catastrophic forgetting within trained models. Graves et al. (2021); Thudi et al. (2022) on the other hand propose retraining the model on corrupted labels for sensitive data providing for a relatively cost-effective measure to induce unlearning. Hessian matrix approximation largely characterizes some of the works (Mehta et al., 2022). Use of fisher information matrices (Becker and Liebig, 2022; Liu et al., 2023b) allows for a direct albeit expensive estimation of the expectation of the hessian due to $\log p(y|x,\theta)$, allowing for strategies pertinent to noising/masking if the said estimation for the data to be forgotten. Fisher forgetting (Becker and Liebig, 2022; Liu et al., 2023b) broadly aims at a similar estimation of the hessians using FIMs in order to conduct informed corruption of samples in the image space. Model decisions on the other hand may lend to some avenues for manipulation (Chen et al., 2023; Wang et al., 2023) in order to induce forgetting using statistical distance based metrics such as $KL$ divergence. Unlearning measures have also experienced a surge in interest within the generative models space and in Large Language Models (LLMs). In the generative model domain (Heng and Soh, 2023a, b; Gandikota et al., 2023), efforts are directed towards inhibiting diffusion or Generative Adversarial Network (GAN)-based models from generating questionable content. Model editing methods (Dai et al., 2021; Geva et al., 2020; Meng et al., 2022) aim to edit factual information encoded within pre-trained LLMs. However, these approaches predominantly focus on selective influential aspects of the trained model in isolation, overlooking the potential for incorporating separate networks to expedite the unlearning process in the model under consideration. Furthermore, in order to incorporate expensive measures for estimation of second order information, significant approximations are formulated which determine the efficiency and performances of the methods.
2.2 Membership Inference Attacks (MIA)
The purpose of a MIA is to uncover from a trained model (target), information revealing whether a sample was involved in the training process of the target model. Assuming black box conditions, the attack model trains itself to make binary predictions using the outputs of the target, inferring belonging or lack thereof of samples in the training set of target. This formulation was first proposed by Shokri et al. (2017), wherein they create multiple shadow models to replicate the behavior of a target model. The shadow models are then utilized to generate data for training a binary classifier to determine membership. Building upon this concept, Salem et al. (2018) enhanced the efficacy of the attack by employing a single shadow model. Refinements to the method were achieved through the integration of supplementary discriminative information (Nasr et al., 2018), which included considerations of labels. Additional contributions (Leino and Fredrikson, 2020; Yeom et al., 2018; Song et al., 2019; Song and Mittal, 2021) introduced evaluative metrics where the membership status of a record is directly ascertained by comparing predefined thresholds based on metrics such as prediction confidences, entropy, or modified entropy of the record. Song and Mittal demonstrated that, by establishing a class-dependent threshold, the metric-based classifier could achieve comparable or even superior inference performance compared to the neural network-based classifier (Song and Mittal, 2021). In this work we adopt the recent work of Yuan and Zhang (2022) towards proposing an effective mechanism for MIA on pruned models. Briefly, the work proposes a transformer based architecture that incorporates multiple sources of information about the target model’s inferential patterns in order to effectively conduct MIA against the model.
Leveraging the advancements in these fields we propose a union of the two, formulating an adversarial game between the defender network, tasked with orchestrating the unlearning process and the attacker network leveraging MIA to infer the membership status of the data to be unlearned. Therefore, under the context of this adversarial framework, the successful expunging of a data point’s influence on the model is indicated when the attacker is unable to distinguish whether the data point originated from the training set or is a representative instance of unseen data.
3 The Proposed Method
Let $\mathcal{D}$ represent the set of training samples utilized for training a network with parameters $\theta_{0}$. The sets $\mathcal{D}_{v}$ and $\mathcal{D}_{t}$ denote the validation and test sets, respectively, employed for evaluating the model, comprising samples that were not part of the training process. $\mathcal{D}_{f}\subseteq\mathcal{D}$ is defined as the subset of samples targeted for removal from the model, resulting in updated parameters $\theta_{u}$. The set $\mathcal{D}_{r}$ constitutes the retained samples, defined as $\mathcal{D}_{r}=\mathcal{D}\setminus\mathcal{D}_{f}$. Two primary scenarios of forgetting are considered: random forgetting, where the model is required to forget samples randomly selected from $\mathcal{D}$, and class-wise forgetting, where samples specific to a certain class are targeted for removal. The benchmark for evaluating an unlearning algorithm is the model trained solely on the retained set. Specific to our method, we introduce in our framework, an attacker model $\mathbf{A}$ trained to infer the membership of a given sample.
From the standpoint of $\mathbf{A}$, in order for a trained defender $\mathbf{D}$ to forget its training due to $\mathcal{D}_{f}$, the attacker must be unable to distinguish between its predictions for $\mathcal{D}_{f}$ and $\mathcal{D}_{v}$. This motivates the algorithm of unlearning wherein $\mathbf{A}$ and $\mathbf{D}$ are trained adversarially with their objectives pit against each other, essentially approaching an optimal convergence, wherein the performance of the gold standard method of training from scratch on $\mathcal{D}_{r}$ is recovered by the defender and the attacker predictions equal $0.5$ for all samples arising from $\mathcal{D}_{f}$ and $\mathcal{D}_{v}$. In what follows, we delve into the specifics of the attacker and defender architectures and objectives and describe our formulation of the problem that allows us to leverage a these networks in the min-max framework, in order to enable unlearning. Following this, we elucidate the integration of the self-supervised objective, enhancing the method’s efficacy and contributing to the overall objective that defines the algorithm.
3.1 The Attacker
Consider a target defender model $\mathbf{D}$, which is trained on the training set $\mathcal{D}$. The objective of the attacker model $\mathbf{A}$, is to deduce a binary prediction indicating the presence or absence of a specific sample in the training set. Typically, the scenario is characterized as a black-box setting, where the attacker’s access is constrained to the outputs of the defender model. The attacker model is trained by utilizing the defender’s inferences on samples originating from both the set of forgotten samples $\mathcal{D}_{f}$ and the validation set $\mathcal{D}_{v}$, with the distinction that the former was used in the training process of the defender, while the latter was not. As we shall see later, it is strategically advantageous to employ a potent attacker that assimilates substantial information to achieve accurate inferences regarding membership. Accordingly, we consider the recent formulation proposed by Yuan and Zhang (2022). Briefly, this attacker is designed to process three crucial pieces of information for each sample:
  • The output predictions tensor generated by the defender for the given sample.
  • The sensitivity of the defender to noisy inputs, evaluated as $\Delta(x)=\frac{1}{n}\sum_{i=1}^{n}\left\|\mathbf{D}(x)-\mathbf{D}(x+\mathbf{\epsilon}_{i})\right\|$ where $\epsilon\sim\mathcal{N}(0,1)$. We evaluate the sensitivity tensor upon averaging over $n=10$ noisy samples.
  • The one-hot label information $y$ of the sample, indicating the class information.
The attacker then formulates the informed inputs upon concatenating these tensors $\mathbf{D}(x_{f})\prime=[\mathbf{D}(x_{f}),\Delta(x_{f}),y_{f}]$, $\mathbf{D}(x_{v})\prime=[\mathbf{D}(x_{v}),\Delta(x_{v}),y_{v}]$, and further encodes using a fully connected (FC) layer. The resulting representation is then propagated through a transformer architecture comprising self-attention layers eventually allowing for a binary prediction score. Through this process, the attacker aims to tease out the influence of the forget set $\mathcal{D}_{f}$ within the defender model and accordingly maximizes the log-likelihood given by $\mathcal{L}_{\mathbf{A}}=\mathbb{E}_{x_{f}\sim\mathcal{D}_{f},x_{v}\sim\mathcal{D}_{v}}[\log(\mathbf{A}(\mathbf{D}(x_{f})\prime)+\log(1-\mathbf{A}(\mathbf{D}(x_{v})\prime)]$.
3.2 The Defender
A network trained on the dataset $\mathcal{D}$ is now tasked with retaining its performance on $\mathcal{D}_{v}$ and $\mathcal{D}_{t}$ whilst forgetting its training on $\mathcal{D}_{f}$. The former may be accomplished using the supervised classification objective wherein the network is provided with labeled samples originating from the $\mathcal{D}_{r}$. This is typically implemented as the cross entropy loss $\mathcal{L}_{\mathbf{D}}=\mathbb{E}_{x_{r}\sim\mathcal{D}_{r}}[-\sum_{i=1}^{K}y_{ri}\cdot\log(\mathbf{D}(x_{r})_{i})]$ for a $K$ class classification problem, which $\mathbf{D}$ aims to minimize. For the latter however, the defender may leverage the capacity of the attacker network to distinguish between its outputs from $\mathcal{D}_{f}$ and $\mathcal{D}_{v}$. Accordingly, in order to effectively forget, the defender network also minimizes the objective of the attacker, therefore defending against an active MIA conducted by the attacker.
3.3 Adversarial Unlearning
The attacker and defender networks are now entangled in adversarial objectives, wherein the defender network not only seeks to defend against the attacker, but also aims to preserve its performance on the validation set $\mathcal{D}_{v}$ and the test set $\mathcal{D}_{t}$ while the attacker is tasked with conducting MIA against the defender. We may therefore formalize the interaction between the attacker and the defender as a min-max game, defined as follows:
$\min_{\mathbf{D}}\max_{\mathbf{A}}\mathcal{V}_{mm}(\mathbf{A},$ $\mathbf{D})\triangleq\mathcal{L}_{\mathbf{D}}+\alpha\mathcal{L}_{\mathbf{A}}$
$=\mathbb{E}_{x_{r}\sim\mathcal{D}_{r}}[-\sum_{i=1}^{K}y_{ri}\cdot\log(\mathbf{D}(x_{r})_{i})]$
$+\alpha[\mathbb{E}_{x_{f}\sim\mathcal{D}_{f},x_{v}\sim\mathcal{D}_{v}}[\log(\mathbf{A}(\mathbf{D}(x_{f})\prime)$
$+\log(1-\mathbf{A}(\mathbf{D}(x_{v})\prime))]],$ (1)
where $\alpha$ modulates the influence of the attacker’s objective on the defender network which also optimizes for a $K$ class classification problem. To facilitate the training of the networks with the specified objective, we employ the iterative updating scheme common to the min-max optimization frameworks for deep network learning (Goodfellow et al., 2020; Arjovsky et al., 2017; Gulrajani et al., 2017), for both the defender and the attacker. The trained defender is provided with samples from $\mathcal{D}_{f}$ and $\mathcal{D}_{v}$ on which it conducts inference. The attacker subsequently uses these inferences to generate its inputs, as outlined in Section 3.1. During this process, the attacker acquires knowledge of the inference patterns exhibited by the defender and endeavors to distinguish between the defender’s outputs for samples from $\mathcal{D}_{f}$ and $\mathcal{D}_{v}$, by maximizing a binary classification objective. This discrimination task is relatively easier for the attacker in the early stages of the algorithm. As the training progresses, the defender becomes aware of the attacker’s motives and consequently adjusts its own learning objectives to defend against the attack, minimizing the attacker’s objective, leading to a decline in its performance on the set of retained samples $\mathcal{D}_{r}$. In response to this degradation, the defender optimizes for learning to classify on $\mathcal{D}_{r}$ to preserve its general capacity for classification on $\mathcal{D}$. This adaptive learning process reflects the dynamic interplay between the defender and the attacker during training, ultimately influencing the defender’s strategies to balance classification performance on the retained set and resistance against the MIA.
One may however consider making practical adjustments to the objective, in consideration of the problem of weaker gradients as identified in the work of Generative Adversatial Nets (GAN) Goodfellow et al. (2020) and instead maximize $\mathbb{E}_{x_{f}\sim\mathcal{D}_{f},x_{v}\sim\mathcal{D}_{v}}[\log(\mathbf{A}(1-\mathbf{D}(x_{f}))+\log(\mathbf{A}(\mathbf{D}(x_{v})))]$ to empirically adjust for the differential capacities of the attacker and the defender networks to counter each other’s efforts. Additionally, one may also incorporate improvements over the GAN algorithm (Arjovsky et al., 2017; Gulrajani et al., 2017; Kodali et al., 2017) in order to ensure better convergence. Incorporation of objectives pertinent to unlearning such as gradient ascent using a negative of the loss function for $\mathcal{D}_{f}$ (Graves et al., 2021; Thudi et al., 2022), or fisher forgetting through additive gaussian noise (Becker and Liebig, 2022; Izzo et al., 2021) is also feasible through relatively nominal adjustments to the algorithm.
3.4 Self-supervised Regularization
The min-max objective strives to render the outputs of $\mathbf{D}$ indistinguishable to a proficient attacker within the output space. Nevertheless, the defender network tends to overcompensate by emphasizing exclusively on the final output layer, aiming to enhance its competence against the attacker, and the feature space discrepancies between $\mathcal{D}_{f}$ and $\mathcal{D}_{v}$ remain overlooked. In order to address these discrepancies we introduce a self supervised regularization term that enables the defender to inherit the task of forgetting samples within the network’s parameters and eases the burden on the final layer of the network. We motivate this objective through the notion that the features generated for the forget set $\mathcal{D}_{f}$ must on average be similar to those arising from the validation set $\mathcal{D}_{v}$ and therefore the test set $\mathcal{D}_{t}$. In our application we consider the feature space redundancy reduction principle utilized by Zbontar et al. (2021) that aims to enforce invariance between features of samples that in principle must be similar. The defender $\mathbf{D}$ essentially consisting of the encoder and the FC layer $\mathbf{D}(.)=\mathbf{D}_{fc}(\mathbf{D}_{enc}(.))$, upon processing the input samples, outputs mean-centered features $\{z_{b}:z_{b}=\mathbf{D}_{enc}(x_{b}),z_{b}\in\mathbb{R}^{D}\}$, where $b$ indexes the batch. Subsequently, for a cross-correlation matrix $\mathcal{C}$ computed between features from $\mathcal{D}_{f}$ and $\mathcal{D}_{v}$, defined as,
$\mathcal{C}_{ij}\triangleq\frac{\sum_{b}z_{b,i}^{\mathcal{D}_{f}}z_{b,j}^{\mathcal{D}_{v}}}{\sqrt{\sum_{b}(z_{b,i}^{\mathcal{D}_{f}})^{2}}\sqrt{\sum_{b}(z_{b,j}^{\mathcal{D}_{t}})^{2}}}$
the self supervised regularization is formulated as follows:
$\min_{D}\mathcal{V}_{ss}(\mathbf{D})\triangleq\sum_{i}(1-\mathcal{C}_{ii})^{2}+\lambda\sum_{i}\sum_{j\neq i}\mathcal{C}_{ij}^{2},$ (2)
where $\lambda$ was set to $-35e\text{-}$ 3 as recommended in Zbontar et al. (2021). The regularization plays a crucial role in augmenting the overall performance of the algorithm, as elucidated through ablation studies in Appendix Section 4.5. However, one may surmise several other variants of the self-supervised objective (Bardes et al., 2021; Caron et al., 2021). Additionally, contrastive objectives (Chen et al., 2020; Sharma et al., 2024; Dwibedi et al., 2021) that enable similarity between $\mathcal{D}_{f}$ and $\mathcal{D}_{v}$ and distance between $\mathcal{D}_{f}$ and $\mathcal{D}_{r}$, may also be incorporated as regularizers. One may also leverage class label information to motivate self-supervised objectives within the classes that the samples belong to. Furthermore, one may incorporate a separate projector network over the defender in order to implement this regularization as is commonly practiced in SSL literature. The overall objective that we utilize in order to train the framework through gradient descent is formulated as:
$\min_{\mathbf{D}}\max_{\mathbf{A}}\mathcal{V}(\mathbf{A},\mathbf{D})\triangleq\mathcal{V}_{mm}(\mathbf{A},\mathbf{D})+\beta\mathcal{V}_{ss}(\mathbf{D}),$ (3)
which concludes our method. Here, $\beta$ signifies the strength of regularization which is empirically determined. The complete procedure for this approach is listed Algorithm 1, which delineates the iterative update schema used to implement the min-max approach, sequentially updating the attacker and the defender networks, leading to the unlearned defender network $\mathbf{D}_{\theta_{u}}$. An illustration of the procedure is provided under Figure 1.
Algorithm 1 Adversarial Unlearning
1:  Input: $\mathcal{D}_{f}$, $\mathcal{D}_{v}$, $\mathcal{D}_{r}$, $\mathbf{A}_{\theta_{\mathbf{A}}}$, $\mathbf{D}_{\theta_{\mathbf{D}}}$
2:  Parameters: Batch Size $B$, Learning Rates $\eta_{\mathbf{A}}$, $\eta_{\mathbf{D}}$, Parameters $\alpha,\beta,\lambda$
3:  for epoch do
4:     Sample $(x_{r},y_{r})\sim\mathcal{D}_{r}$
5:     Evaluate $\mathbf{D}(x_{f})\prime\,\forall x_{f}\in\mathcal{D}_{f}$, $\mathbf{D}(x_{v})\prime\,\forall x_{v}\in\mathcal{D}_{v}$
6:     Evaluate $\mathcal{V}_{ss}\,\forall(x_{f},x_{v})\in(\mathcal{D}_{f},\mathcal{D}_{v})$
7:     Update attacker by stochastic gradient ascent as
$\theta_{\mathbf{A}}\leftarrow\theta_{\mathbf{A}}+\eta_{\mathbf{A}}\triangledown_{\mathbf{\theta_{\mathbf{A}}}}\mathcal{L}_{\mathbf{A}},\text{ with }\mathcal{L}_{\mathbf{A}}=\frac{1}{B}\sum\log(\mathbf{A}(\mathbf{D}(x_{f})\prime)+\log(1-\mathbf{A}(\mathbf{D}(x_{v})\prime)$
8:     Update defender by stochastic gradient descent as
$\theta_{\mathbf{D}}\leftarrow\theta_{\mathbf{D}}-\eta_{\mathbf{D}}\triangledown_{\mathbf{\theta_{\mathbf{D}}}}(\mathcal{L}_{\mathbf{D}}+\alpha\mathcal{L}_{\mathbf{A}}+\beta\mathcal{V}_{ss}),\text{ with }\mathcal{L}_{\mathbf{D}}=\frac{1}{B}\sum\Big{[}-\sum_{i=1}^{K}y_{ri}\cdot\log(\mathbf{D}(x_{r})_{i})\Big{]}$
9:  end for
Table 1: Cifar-10, Cifar-100 comparison. We evaluate the unlearning performance on Cifar datasets. The top two tables encompass the unlearning performances on Cifar-10 for the methods listed for random and class-wise schemes respectively. Evaluation metrics from section 4 are employed, including the averaged disparity of performances relative to retraining on the retain set $\mathcal{D}r$. Five separate seeds are utilized for each method, and the reported mean and standard deviation are presented in the format $a{\pm b}$. Additionally, unlearning is performed on $95\%$ sparse models for each method, employing OMP (Ma et al., 2021) to enforce sparsity. The relative performances are compared against the ”retrain” standard, and the average disparity provides an overarching assessment of all algorithms across the employed metrics. (UA: Unlearning Acc., RA: Retaining Acc., TA: Test Acc.)
Cifar-10 Random Forgetting Methods UA MIA-Efficacy RA TA Avg. Disparity Run Time Dense 95% Sparse Dense 95% Sparse Dense 95% Sparse Dense 95% Sparse Dense 95% Sparse min. Retrain $5.80_{\pm 0.12}$ $6.81_{\pm 0.23}$ $13.91_{\pm 0.15}$ $15.17_{\pm 0.15}$ $100.00_{\pm 0.00}$ $100.00_{\pm 0.00}$ $94.30_{\pm 0.13}$ $92.21_{\pm 0.22}$ 0.00 0.00 82.15 FT $0.18_{\pm 0.04}$ $0.04_{\pm 0.04}$ $1.70_{\pm 0.10}$ $1.95_{\pm 0.01}$ $99.92_{\pm 0.07}$ $99.98_{\pm 0.01}$ $94.25_{\pm 0.15}$ $94.13_{\pm 0.11}$ $4.48$ $4.52$ 4.91 GA $0.00_{\pm 0.00}$ $0.03_{\pm 0.01}$ $0.35_{\pm 0.15}$ $0.69_{0.03}$ $99.99_{\pm 0.01}$ $100.00_{\pm 0.00}$ $94.80_{\pm 0.03}$ $94.40_{\pm 0.04}$ 4.96 5.86 0.37 FF $6.96_{\pm 1.15}$ $10.11_{\pm 0.48}$ $10.24_{\pm 0.52}$ $10.47_{\pm 0.48}$ $93.09_{\pm 1.06}$ $90.02_{\pm 0.05}$ $88.51_{\pm 0.98}$ $85.13_{\pm 0.01}$ 4.37 6.26 42.9 IU $0.42_{\pm 0.78}$ $0.48_{\pm 0.36}$ $1.49_{\pm 1.18}$ $2.58_{\pm 0.75}$ $99.57_{\pm 0.81}$ $99.54_{\pm 0.37}$ $93.89_{\pm 1.03}$ $92.97_{\pm 0.55}$ 4.96 5.95 4.93 Ours $3.46_{\pm 0.13}$ $4.92_{\pm 0.26}$ $8.25_{\pm 0.02}$ $12.27_{\pm 0.00}$ $99.50_{\pm 0.21}$ $96.78_{\pm 0.28}$ $93.50_{\pm 0.12}$ $90.96_{\pm 0.24}$ $2.32$ $2.32$ 7.98 Cifar-10 Class-wise Forgetting Methods UA MIA-Efficacy RA TA Avg. Disparity Run Time Dense 95% Sparse Dense 95% Sparse Dense 95% Sparse Dense 95% Sparse Dense 95% Sparse min. Retrain $100.00_{\pm 0.00}$ $100.00_{\pm 0.00}$ $100.00_{\pm 0.00}$ $100.00_{\pm 0.00}$ $100.00_{\pm 0.00}$ $100.00_{\pm 0.00}$ $94.81_{\pm 0.09}$ $91.76_{\pm 0.93}$ 0.00 0.00 82.00 FT $8.36_{\pm 3.03}$ $31.31_{\pm 17.01}$ $40.76_{\pm 8.03}$ $75.29_{\pm 16.94}$ $99.92_{\pm 0.03}$ $99.98_{\pm 0.01}$ $94.41_{\pm 0.29}$ $94.48_{\pm 0.26}$ 37.84 24.03 4.88 GA $93.80_{\pm 1.96}$ $96.69_{\pm 1.97}$ $96.06_{\pm 1.87}$ $97.37_{\pm 2.24}$ $93.76_{\pm 0.73}$ $88.01_{\pm 2.16}$ $87.15_{\pm 0.58}$ $82.29_{\pm 1.98}$ 6.00 6.84 0.39 FF $67.57_{\pm 13.05}$ $92.54_{\pm 3.59}$ $100.00_{\pm 0.00}$ $100.00_{\pm 0.0}$ $100.00_{\pm 0.00}$ $100.00_{\pm 0.00}$ $94.98_{\pm 0.21}$ $94.75_{\pm 0.29}$ 8.14 3.42 42.7 IU $64.41_{\pm 24.94}$ $95.60_{\pm 6.15}$ $81.66_{\pm 17.64}$ $98.95_{\pm 1.81}$ $98.88_{\pm 1.33}$ $97.99_{\pm 1.47}$ $92.77_{\pm 1.75}$ $91.47_{\pm 1.79}$ 14.27 1.94 4.40 Ours $100.00_{\pm 0.00}$ $100.00_{\pm 0.00}$ $100.00_{\pm 0.00}$ $100.00_{\pm 0.00}$ $98.90_{\pm 0.20}$ $95.00_{\pm 0.43}$ $93.22_{\pm 0.74}$ $90.47_{\pm 0.48}$ 0.67 1.57 9.91
Cifar-100 Random Forgetting Methods UA MIA-Efficacy RA TA Avg. Disparity Run Time Dense 95% Sparse Dense 95% Sparse Dense 95% Sparse Dense 95% Sparse Dense 95% Sparse min. Retrain $24.75_{\pm 0.11}$ $26.93_{\pm 1.04}$ $49.68_{\pm 0.35}$ $44.49_{\pm 0.44}$ $99.98_{\pm 0.01}$ $99.57_{\pm 0.08}$ $74.57_{\pm 0.06}$ $69.73_{\pm 0.15}$ 0.00 0.00 82.33 FT $0.11_{\pm 0.03}$ $2.57_{\pm 0.21}$ $5.66_{\pm 0.47}$ $12.96_{\pm 0.25}$ $99.97_{\pm 0.01}$ $99.57_{\pm 0.04}$ $75.45_{\pm 0.17}$ $73.55_{\pm 0.17}$ 16.65 13.02 4.91 GA $0.04_{\pm 0.01}$ $0.21_{\pm 0.05}$ $2.07_{\pm 0.15}$ $4.26_{\pm 0.29}$ $99.98_{\pm 0.00}$ $99.83_{\pm 0.02}$ $75.38_{\pm 0.07}$ $74.33_{\pm 0.08}$ 17.58 15.52 0.41 FF $0.05_{\pm 0.01}$ $0.23_{\pm 0.06}$ $2.04_{\pm 0.21}$ $4.12_{\pm 0.15}$ $99.98_{\pm 0.00}$ $99.84_{\pm 0.01}$ $75.48_{\pm 0.07}$ $74.26_{\pm 0.05}$ 17.56 15.57 42.8 IU $0.11_{\pm 0.06}$ $2.49_{\pm 0.91}$ $3.65_{\pm 0.70}$ $7.43_{\pm 0.89}$ $99.94_{\pm 0.04}$ $98.07_{\pm 0.81}$ $74.72_{\pm 0.34}$ $71.83_{\pm 0.75}$ 17.35 15.22 5.25 Ours $10.50_{\pm 0.45}$ $13.36_{\pm 0.97}$ $30.09_{\pm 0.01}$ $24.65_{\pm 0.01}$ $97.60_{\pm 0.45}$ $95.21_{\pm 0.54}$ $71.25_{\pm 0.59}$ $69.94_{\pm 0.01}$ 5.48 5.43 13.3 Cifar-100 Class-wise Forgetting Methods UA MIA-Efficacy RA TA Avg. Disparity Run Time Dense 95% Sparse Dense 95% Sparse Dense 95% Sparse Dense 95% Sparse Dense 95% Sparse min. Retrain $100.00_{\pm 0.00}$ $100.00_{\pm 0.00}$ $100.00_{\pm 0.00}$ $100.00_{\pm 0.00}$ $99.98_{\pm 0.01}$ $96.59_{\pm 0.12}$ $73.75_{\pm 0.20}$ $69.49_{\pm 0.21}$ 0.00 0.00 82.22 FT $12.04_{\pm 7.1}$ $59.15_{\pm 23.02}$ $70.22_{\pm 17.57}$ $78.67_{\pm 19.74}$ $99.95_{\pm 0.01}$ $98.43_{\pm 0.09}$ $74.60_{\pm 0.30}$ $72.44_{\pm 0.32}$ 29.23 14.35 4.85 GA $68.67_{\pm 1.16}$ $84.22_{\pm 4.19}$ $91.87_{\pm 6.36}$ $91.56_{\pm 5.69}$ $96.55_{\pm 3.02}$ $89.91_{\pm 6.20}$ $68.40_{\pm 3.42}$ $65.85_{\pm 4.13}$ 12.06 8.64 0.38 FF $0.05_{\pm 0.09}$ $0.17_{\pm 0.26}$ $1.99_{\pm 1.13}$ $4.27_{\pm 3.03}$ $99.98_{\pm 0.00}$ $99.83_{\pm 0.00}$ $75.43_{\pm 0.02}$ $74.34_{\pm 0.00}$ 49.07 46.87 42.6 IU $20.14_{\pm 15.76}$ $84.09_{\pm 21.58}$ $82.67_{\pm 9.40}$ $95.24_{\pm 7.49}$ $99.97_{\pm 0.02}$ $96.42_{\pm 2.68}$ $74.69_{\pm 0.46}$ $70.44_{\pm 2.24}$ 24.67 4.97 4.93 Ours $97.20_{\pm 2.15}$ $99.27_{\pm 0.76}$ $99.84_{\pm 0.02}$ $99.93_{\pm 0.01}$ $98.78{\pm 0.26}$ $91.52{\pm 1.06}$ $71.49{\pm 0.52}$ $69.60{\pm 0.63}$ 0.89 0.85 16.65
4 Experiments
4.1 Criteria
In an ideal scenario for machine unlearning, a perfectly functional system ensures the following set of conditions are met.
  • The model must be able to generalize to other datasets. We evaluate this condition through the accuracy of the scrubbed defender on the test set $\mathcal{D}_{t}$ Formally, $\mathbf{TA}=Acc_{\mathcal{D}_{t}}(\mathbf{D}_{\theta_{u}})$.
  • The performance of the unlearned model is retained on the $\mathcal{D}_{r}$ which is measured through retain set accuracy. Formally, $\mathbf{RA}=Acc_{\mathcal{D}_{r}}(\mathbf{D}_{\theta_{u}})$.
  • The unlearning performance of the model closely approximates the gold standard of retraining, which is measured on the forget set $\mathcal{D}_{f}$ using unlearning accuracy defined as $\mathbf{UA}=1-Acc_{\mathcal{D}_{f}}(\mathbf{D}_{\theta_{u}})$.
  • The general capacity of the unlearned model to defend against MIA. For this metric we employ a separate prediction confidence based attack, which utilizes the differences in confidence of the trained network for samples arising from $\mathcal{D}$ vs $\mathcal{D}_{t}$ or $\mathcal{D}_{v}$. The confidence based predictor (Yeom et al., 2018; Song et al., 2019) is formulated through training a support vector machine on the prediction confidences of the network for samples arising from $\mathcal{D}$ and $\mathcal{D}_{t}$ and subsequently used to predict the origin of $\mathcal{D}_{f}$. A successful attack will therefore perform with high MIA-Efficacy on $\mathcal{D}_{f}$. Formally, for an attacker to fail, it must predict samples arising from $\mathcal{D}_{f}$ as true negatives and therefore, $\mathbf{MIA\text{-}efficacy}=TN/|\mathcal{D}_{f}|$ where $TN$ denotes the true negative inferences by the attacker.
  • The efficiency of the algorithm is evaluated by considering the run-time for one random seed of forgetting, ensuring it is not overly cumbersome compared to the gold standard retraining model and other unlearning methods.
Our comparisons are set against popular paradigms of machine unlearning described as follows.
  • Fine-tuning (FT) (Warnecke et al., 2021; Golatkar et al., 2020): This method involves finetuning the network only on the retain set $\mathcal{D}_{r}$ in order to induce amnesia over the network’s prior training on $\mathcal{D}_{f}$ leveraging the network’s tendency for catastrophic forgetting, and ensconce the performance over the retain set.
  • Gradient-ascent (GA) (Graves et al., 2021; Thudi et al., 2022): Using the negative of the classification loss evaluated on $\mathcal{D}_{f}$, this method retrains the model in order to explicitly negate the prior training and induce forgetting.
  • Fisher Forget (FF) (Becker and Liebig, 2022; Izzo et al., 2021): FF uses estimates the hessian of the defender network using fisher information matrix, in order to informatively corrupt the of the pre-trained network’s parameters $\theta_{0}$ using additive gaussian noise in the image space.
  • Influence Unlearning (IU) (Koh and Liang, 2017): This method uses estimates of influence functions which evaluate the change in parameters of a pre-trained network due to the presence/absence of samples in $\mathcal{D}$ through ”upweighing” (Cook and Weisberg, 1980) these samples on the network parameters, allowing for the method to manipulate the influence of the loss due to $\mathcal{D}_{f}$ on the defender.
4.2 Configuration
Training is conducted on a single Nvidia RTX-A6000 GPU. The defender’s base model is the ResNet-18 (He et al., 2016), while the attacker network adopts a transformer structure with $3$ multihead self-attention layers and $4$ attention heads, retaining the default configuration from Yuan and Zhang (2022). Pre-training of the defender employs a learning rate of $0.1$ for both Cifar datasets. For random forgetting, the (un-)learning rate is set to $0.01$ with the dense model, while a higher rate of $0.03$ is found to be more suitable for a sparse model in this scenario. In the class-wise unlearning scenario, a rate of $0.02$ is used for both levels of sparsity and datasets. Empirical adjustments are made to the $\alpha$ and $\beta$ parameters, setting them to $0.9$ and $0.001$, respectively.
Training Adjustments
We find it beneficial to pre-train the attacker network for $1000$ iterations to classify the defender’s outputs for $\mathcal{D}$ and $\mathcal{D}_{v}$, enabling the framework to leverage informative gradients provided by the attacker in the early training stages. During class-wise forgetting, we observe that unlearning the class for both datasets is notably easier than retaining general performance. Consequently, we adjust the $\alpha$ parameter to 0 after the first 30 iterations in this scenario.
Figure 2: MIA-Attack Robustness. We incorporate additional MIA-Attack methods and compare the robustness of our method against the baselines. The attacks are based on Prediction Confidence, Prediction Entropy, Modified Prediction Entropy and Prediction Probability as elucidated in Section 4.4 under Appendix.
4.3 Results
Our results are listed in the Table 1 for all the aforementioned configurations. We conduct our evaluations under the random forgetting scheme where $10\%$ of the samples are chosen randomly to be scrubbed from the pre-trained model, and class-wise forgetting scheme where all samples from a randomly chosen class are tasked for removal. All methods were evaluated using the dense model and a $95\%$ sparse model, in order to gauge the robustness and efficacy of the listed methods under varying conditions of sparsity. Sparsity in the networks is enforced using One-Shot Magnitude Pruning (OMP) (Ma et al., 2021) due to its lower computational overhead and better generalization (Jia et al., 2023). The gold-standard for unlearning which is to retrain the model on $\mathcal{D}_{r}$ is also evaluated. The assessment employs the metrics outlined in section 4.1. It is imperative to emphasize that elevated values across these metrics do not inherently denote superior performance and unlearning methods should exhibit a minimal disparity in comparison to the retrain method. Consequently, our evaluation incorporates the computation of the average disparity of scores, aiming to simulate a measure of proximity with this gold standard. The experiments reveal that our proposed method surpasses all baseline approaches across the utilized metrics and attains the utmost proximity to the gold standard. In the context of the random forgetting scheme, our method consistently achieves the closest performance to the retrain method. Simultaneously, it maintains superior performance in both the retain set (RA) and test set (TA). Contrastingly, other baseline methods exhibit ineffectiveness in the task of forgetting, resulting in excessively high scores for RA and TA, or they adopt over-compensatory measures, causing a decline in the model’s overall classification capacity,, as evidenced by the MIA-efficacy metric. Notably, FF achieves significant scores in Cifar-10 under this scheme but experiences a considerable decline with over $5\%$ point gap from the retrain benchmark in the TA and RA metrics.
Nevertheless, notable observations arise under the class-wise scheme for both datasets, where the methods demonstrate competitive performances. In the case of the Cifar-10 dataset, a high UA metric is observed across the methods, accompanied by elevated values for RA and TA. Notably, our method achieves flawless performance across the UA and MIA-Efficacy metrics, maintaining a remarkably narrow margin for the RA and TA metrics with merely a $0.67$ score in the Avg. Disparity measure. This achievement mitigates the necessity for sparsification measures to improve performance. For Cifar-100, a comparable performance is also observed across these metrics, while concurrently sustaining a minimal performance gap with the gold standard.
Sparse models however, ease the process of unlearning performance for the methods as noted by Jia et al. (2023) whilst incurring a marginal cost over the TA and RA metrics. This phenomenon is particularly pronounced in class-wise forgetting schemes, where the baseline methods demonstrate significant improvements across the metrics and resulting in substantially lower Avg. Disparity scores. Our approach demonstrates a comparable trend, closely aligning with the retraining method across various metrics. Notably, we argue that incorporating sparsity as a performance enhancement measure yields marginal benefits in our context. The algorithm consistently achieves robust performance with dense variants, presenting challenges for the baselines, even when employing $95\%$ sparse models. Moreover, $\mathcal{V}_{ss}$ plays a critical role in enhancing performance as we elucidate in Section 4.5.
Table 2: Ablations for $\mathcal{V}_{ss}$ Here we evaluate our method on the random forgetting scenario over both Cifar-10 and Cifar-100 datasets to illustrate the role of the self-supervised regularization objective in our algorithm. The format follows from Table 1. (UA: Unlearning Acc., RA: Retaining Acc., TA: Test Acc.)
Cifar-10 Random Forgetting Methods UA MIA-Efficacy RA TA Avg. Disparity Run Time Dense 95% Sparse Dense 95% Sparse Dense 95% Sparse Dense 95% Sparse Dense 95% Sparse min. Retrain $5.80_{\pm 0.12}$ $6.81_{\pm 0.23}$ $13.91_{\pm 0.15}$ $15.17_{\pm 0.15}$ $100.00_{\pm 0.00}$ $100.00_{\pm 0.00}$ $94.30_{\pm 0.13}$ $92.21_{\pm 0.22}$ 0.00 0.00 82.15 Ours $3.46_{\pm 0.13}$ $4.92_{\pm 0.26}$ $8.25_{\pm 0.02}$ $12.27_{\pm 0.01}$ $99.50_{\pm 0.21}$ $96.78_{\pm 0.28}$ $93.50_{\pm 0.12}$ $90.96_{\pm 0.24}$ $2.32$ $2.32$ 7.98 Ours (No $\mathcal{V}_{ss}$) $3.03_{\pm 0.23}$ $4.13_{\pm 0.28}$ $7.60_{\pm 0.02}$ $11.67_{\pm 0.02}$ $99.64_{\pm 0.21}$ $97.04_{\pm 0.28}$ $93.54_{\pm 0.19}$ $91.04_{\pm 0.45}$ $2.61$ $2.59$ 7.83 Cifar-100 Random Forgetting Methods UA MIA-Efficacy RA TA Avg. Disparity Run Time Dense 95% Sparse Dense 95% Sparse Dense 95% Sparse Dense 95% Sparse Dense 95% Sparse min. Retrain $24.75_{\pm 0.11}$ $26.93_{\pm 1.04}$ $49.68_{\pm 0.35}$ $44.49_{\pm 0.44}$ $99.98_{\pm 0.01}$ $99.57_{\pm 0.08}$ $74.57_{\pm 0.06}$ $69.73_{\pm 0.15}$ 0.00 0.00 82.33 Ours $10.50_{\pm 0.45}$ $13.36_{\pm 0.97}$ $30.09_{\pm 0.01}$ $24.65_{\pm 0.01}$ $97.60_{\pm 0.45}$ $95.21_{\pm 0.54}$ $71.25_{\pm 0.59}$ $69.94_{\pm 0.01}$ 5.48 5.43 13.3 Ours (No $\mathcal{V}_{ss}$) $9.79_{\pm 0.51}$ $13.34_{\pm 0.97}$ $29.10_{\pm 0.32}$ $23.34_{\pm 0.07}$ $97.62_{\pm 0.41}$ $95.73_{\pm 0.53}$ $71.28_{\pm 0.44}$ $70.62_{\pm 0.05}$ 5.71 5.57 12.3
4.4 MIA-Robustness
We perform Membership Inference Attacks (MIAs) based on metrics on the resulting unlearned model. These attacks involve deducing membership by calculating various metrics on the prediction vectors. Unlike binary classification predictions based on outputs of the defender network, metric based MIA leverage informative metrics in order to elucidate intricate patterns in the defender’s outputs for members and non members. We incorporate four distinct metrics, described as follows.
  • Prediction correctness MIA (Yeom et al., 2018): The attacker is trained to infer membership of a record if the defender’s predictions are correct. This attack capitalizes on the defender’s ability to predict accurately on $\mathcal{D}$ which may not generalize to $\mathcal{D}_{t}$.
  • Prediction confidence MIA (Salem et al., 2018): The attacker infers membership based on the maximum prediction confidence scores of the defender across all samples. This leverages differential scores of confidence measures for the defender over $\mathcal{D}$ and $\mathcal{D}_{t}$.
  • Prediction entropy MIA (Salem et al., 2018): The attacker evaluates the entropy of predictions over samples from $\mathcal{D}$ and $\mathcal{D}_{t}$, considering the possibility that the entropy of predictions for $\mathcal{D}_{t}$ is higher.
  • Modified prediction entropy MIA (Song et al., 2019): In addition to prediction entropy, the attacker incorporates ground truth label information to leverage the defender’s capacity for false negatives, inferring membership based on, for example, incorrectly predicted records.
The attackers are trained separately using SVMs over inferences from $\mathbf{D}_{\theta_{u}}$, The illustration is provided in Figure 2. We notice that the framework having received explicit information solely in the form of prediction scores and ground truth labels manages to generalize to foreign attackers trained independently, consistently exhibiting superior performances across the metrics and yet again closely approximating the retrained model.
4.5 Role of $\mathcal{V}_{ss}$
We conduct ablations in order to evaluate the role of the self supervised objective (Section 3.4) towards the overall goal of effective machine unlearning and robust generalization. The ablations are listed in Table 2. We notice that the incorporation of the objective consistently yields a higher UA and MIA-efficacy, with marginally lower scores under the RA and TA metrics. Henceforth, this objective serves as an optional augmentation lending to overall benefits as measured by Avg. Disparity at the price of higher run-times. Nevertheless, the performance of the method remains robust and one may opt to incorporate the objective within the framework through marginal adjustments using the parameter $\beta$, as per the use case and empirical needs.
5 Conclusion
In this study, we employ a straightforward and differentiable framework, adhering to the well-established principles of min-max optimization, to address the challenge of machine unlearning. This framework demonstrates high efficiency, yielding substantial runtime improvements over retraining (at least 5x in the case of Cifar-100 class-wise). Moreover, its effectiveness is evident through consistently achieving the lowest disparity scores across various configurations. Remarkably, the proposed method attains near-perfect performance, particularly notable in the class-wise forgetting schemes. In these scenarios, the unlearning accuracy and MIA robustness reach 100% for Cifar-10 and approach this level for Cifar-100, even within dense network regimes with minimal compromises observed in the test accuracies. Furthermore, the method avoids the expenses of hessian matrix approximations for deep networks and involves minimal parameter tuning requirements. The approach additionally, is adaptable to incorporation of advanced MIA attack techniques and improvements over the objectives of the defender and attacker in order to ensure better performances and convergence. Nevertheless, it is noteworthy that the framework excels under the class-wise unlearning scheme, leaving room for refinement in the context of the random unlearning scheme—potentially a more realistic scenario for machine unlearning. We hypothesize that integrating improved optimization procedures (Arjovsky et al., 2017; Gulrajani et al., 2017; Kodali et al., 2017) and employing stronger, multiple attackers leveraging additional information from predictions (Yeom et al., 2018; Salem et al., 2018; Song and Mittal, 2021) could enhance the robustness of the framework in this scheme. This avenue remains a subject for future exploration.
6 Broader Impact
This paper on machine unlearning introduces impactful advancements with far-reaching implications. Our proposed techniques not only enhance privacy protection by enabling the removal or modification of sensitive information from trained models but also contribute to fostering fairness by allowing models to adapt and rectify biases over time. Moreover, the efficiency of our machine unlearning methods, in contrast to traditional retraining approaches, aligns with sustainability goals by reducing computational resource requirements. The transparent and accountable nature of our unlearning techniques enhances trust in AI technologies, crucial for widespread acceptance and cooperation between AI systems and users. As researchers and practitioners, we emphasize the responsible navigation of these impacts, ensuring that our machine unlearning strategies adhere to ethical principles and societal values.